# On the kernels of the pro-$p$ outer Galois representations associated to once-punctured CM elliptic curves

### v2

@@ 2.1. Analogue of Ihara’s power series and Soulé characters
$\Psi^{{\dagger}}\coloneqq\{f\in\mathrm{Aut}\left(\Pi_{1,1}/[\Pi_{1,1}(2),\Pi_{1,1}(2)]\right)\mid\text{$f$ preserves $\langle\bar{z}\rangle$}\}$
and
@@
$\Psi^{{\dagger}}_{1}\coloneqq\ker\left(\Psi^{{\dagger}}\to\mathrm{Aut}(\Pi^{\rm{ab}})\right),$
where $\bar{z}\in\Pi_{1,1}(2)/[\Pi_{1,1}(2),\Pi_{1,1}(2)]$ is the image of $z$ under the natural projection. We shall identify $\mathbb{Z}_{p}[[\Pi_{1,1}^{\rm{ab}}]]$ with $\mathbb{Z}_{p}[[T_{1},T_{2}]]$ via $T_{i}\coloneqq x_{i}-1$ for $i=1,2$. The action of $\mathrm{Aut}(\Pi_{1,1}^{{\rm ab}})\cong\mathrm{GL}_{2}(\mathbb{Z}_{p})$ on $\Pi_{1,1}^{{\rm ab}}$ naturally extends to that of $\mathbb{Z}_{p}[[\Pi_{1,1}^{{\rm ab}}]]\cong\mathbb{Z}_{p}[[T_{1},T_{2}]]$.
Every element $f$ of $\Psi^{{\dagger}}_{1}$ is determined by a pair
$f(x_{i})x_{i}^{-1}\in\Pi_{1,1}(2)/[\Pi_{1,1}(2),\Pi_{1,1}(2)]\quad\text{for $i=1,2$}.$
Since $\Pi_{1,1}(2)/[\Pi_{1,1}(2),\Pi_{1,1}(2)]$ is a free $\mathbb{Z}_{p}[[T_{1},T_{2}]]$-module generated by $\bar{z}$ [Iha86, Theorem 2], there exists a unique power series $G_{i}(f)\in\mathbb{Z}_{p}[[T_{1},T_{2}]]$ satisfying
$f(x_{i})x_{i}^{-1}=G_{i}(f)z\quad\text{for $i=1,2$}.$
@@
For every $f\in\Psi^{{\dagger}}_{1}$, two power series $G_{1}(f)$ and $G_{2}(f)$ satisfy
@@
If we write $H(f)\coloneqq\frac{G_{2}(f)}{T_{2}}=\frac{G_{1}(f)}{T_{1}}$, then $H\colon\Psi_{1}^{{\dagger}}\to\mathbb{Z}_{p}[[T_{1},T_{2}]]$ is an isomorphism. Moreover, the $\mathrm{GL}_{2}(\mathbb{Z}_{p})$-action on $\Psi_{1}^{{\dagger}}$ induced by the following exact sequence
$1\to\Psi_{1}^{{\dagger}}\to\Psi\to\mathrm{GL}_{2}(\mathbb{Z}_{p})\to 1$
makes $H\colon\Psi_{1}^{{\dagger}}\xrightarrow{\sim}\mathbb{Z}_{p}[[T_{1},T_{2}]](1)$ equivariant. Here, $(1)$ denotes the twist by the determinant character.
@@
$s\colon G_{K(E[p^{\infty}])}\to F^{3}\Gamma^{{\dagger}}_{1,1}$
@@
Hence the image of $G_{K(E[p^{\infty}])}$ under $\rho_{X,p}$ is contained in the third component $F^{3}\Gamma_{1,1}$. By composing the inverse of the natural projection $F^{3}\Gamma_{1,1}^{{\dagger}}\xrightarrow{\sim}F^{3}\Gamma_{1,1}$, we obtain a homomorphism $s\colon G_{K(E[p^{\infty}])}\to F^{3}\Gamma^{{\dagger}}_{1,1}$ which lifts $\rho_{X,p}\mid_{G_{K(E[p^{\infty}])}}$.
@@
We define a $\mathrm{Gal}(K(E[p^{\infty}])/K)$-equivariant homomorphism
$\alpha_{1,1}\colon G_{K(E[p^{\infty}])}\to\mathbb{Z}_{p}[[T_{1},T_{2}]](1)$
as a composite of
@@
    the isomorphism $H\colon\Psi^{{\dagger}}_{1}\xrightarrow{\sim}\mathbb{Z}_{p}[[T_{1},T_{2}]](1)$ constructed in Lemma 2.2.
In [Nak95], Nakamura obtained the following explicit formula of the power series $\alpha_{1,1}$ for general once-punctured elliptic curves. Note that we use the convention $0^{0}\coloneqq 1$ and regard $\mathbb{Z}_{p}[[T_{1},T_{2}]]$ as a subring of $\mathbb{Q}_{p}[[U_{1},U_{2}]]$ where $U_{i}\coloneqq\log(1+T_{i})$ in the following formula.
@@
holds for every $\sigma\in G_{K(E[p^{\infty}])}$, where $\kappa_{\boldsymbol{m}}\colon G_{K(E[p^{\infty}])}^{\rm{ab}}\to\mathbb{Z}_{p}$ is a Kummer character whose reduction modulo $p^{n}$ corresponds to the $p^{n}$-th roots of
@@
for every $n\geq 1$. Here $\theta(z,L)$ is the fundamental theta function defined in [Nak95, (2.1)], and the elements $(\omega_{1,n})_{n}\in T_{\mathfrak{p}}(E)$ and $(\omega_{2,n})_{n}\in T_{\bar{\mathfrak{p}}}(E)$ are defined as in Lemma 2.1 (1).
Remark 2.4.
Note that $\theta(z,L)$ is defined on $\mathbb{C}$, not on $E(\mathbb{C})$. Hence we have to choose a lift of $a\omega_{1,n}+b\omega_{2,n}\in E(\mathbb{C})\cong\mathbb{C}/L$ to $\mathbb{C}$, to consider the value $\theta(a\omega_{1,n}+b\omega_{2,n},L)$. However, by [Nak95, (2.2) Proposition (1)], the values of $\theta(z,L)$ at any such two lifts are equal up to $p$-th power roots of unity. Since $K(E[p^{\infty}])$ contains all $p$-th power roots of unity, the Kummer characters associated to $p$-th power roots of $\theta(z,L)$ at any such two lifts coincide with each other. This explains the notation in Theorem 2.3 above.
In the case of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$, it is known that coefficients of Ihara’s power series correspond to the Soulé characters $\{\chi_{m}\}_{m\geq 3,\mathrm{odd}}$ (see Ihara-Kaneko-Yukinari [IKY87, Theorem $\mathrm{A}_{2}$], for example). We therefore define their elliptic analogues as follows:
@@
For every $\boldsymbol{m}=(m_{1},m_{2})>\boldsymbol{1}$ such that $\lvert\boldsymbol{m}\rvert$ is even, we call the character $\kappa_{\boldsymbol{m}}$ appearing in Theorem 2.3 the $\boldsymbol{m}$-th elliptic Soulé character associated to $X$.
Nakamura proved that some linear combinations of the elliptic Soulé characters are nontrivial [Nak95, (3.12)]. In [Ish25], we conditionally proved that the elliptic Soulé characters arising from once-punctured CM elliptic curves are nontrivial. We briefly summarize previous results.
Since $\alpha_{1,1}$ is compatible with the action of $\mathrm{Gal}(K(E[p^{\infty}])/K)$, the $\boldsymbol{m}$-th elliptic Soulé character $\kappa_{\boldsymbol{m}}$ defines a $\mathrm{Gal}(K(E[p^{\infty}])/K)$-equivariant homomorphism
$\kappa_{\boldsymbol{m}}\colon G_{K(E[p^{\infty}])}^{\mathrm{ab}}\to\mathbb{Z}_{p}(\boldsymbol{m})$
for every $\boldsymbol{m}=(m_{1},m_{2})>\boldsymbol{1}$ such that $\lvert\boldsymbol{m}\rvert$ is even. Let
$I\coloneqq\{\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 1}^{2}\setminus\{\boldsymbol{1}\}\mid m_{1}\equiv m_{2}\bmod\lvert O_{K}^{\times}\rvert\},$
and recall that $\Omega$ is the maximal pro-$p$ extension of $K(p)$ unramified outside $p$.
@@
The $\boldsymbol{m}$-th elliptic Soulé character $\kappa_{\boldsymbol{m}}$ is trivial unless $\boldsymbol{m}$ is contained in $I$. For $\boldsymbol{m}\in I$, the character $\kappa_{\boldsymbol{m}}$ factors through $\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}$ and, moreover, is nontrivial if $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(\boldsymbol{m}))$ is finite.
Here $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(\boldsymbol{m}))$ denotes the second étale cohomology group of the spectrum of the ring of $p$-integers of $K$ with coefficient in $\mathbb{Z}_{p}(\boldsymbol{m})$, whose finiteness is a special case of a conjecture of Jannsen [Jan89, Conjecture 1]. For example, such a cohomology group is known to vanish for every $\boldsymbol{m}\in(p-1)\mathbb{Z}_{\geq 1}^{2}$ by e.g. [Ish25, Lemma 4.7]. It is also finite, when the coefficient is the $m$-th Tate twist $\mathbb{Z}_{p}(m)$ for some $m\geq 2$ [Sou79, page 287, Corollaire].
The following result gives a sufficient condition that certain elliptic Soulé characters are surjective:
@@
    If the class number of $K(p)$ is not divisible by $p$ and there are only two primes of $K(p)$ above $p$, then $\kappa_{\boldsymbol{m}}$ is surjective for every $\boldsymbol{m}\in I$ such that $\boldsymbol{m}\not\equiv\boldsymbol{1}\bmod p-1$.
Although $\kappa_{\boldsymbol{m}}$ is not surjective for every $\boldsymbol{m}\in I$ such that $\boldsymbol{m}\geq(2,2)$ and $\boldsymbol{m}\equiv\boldsymbol{1}\mod p-1$, we will see that it is nontrivial under the assumption of Theorem A, cf. Corollary A.3 in the appendix.
2.2. Analogue of the Deligne-Ihara conjecture and its consequence
In this subsection, we introduce a certain filtration on the absolute Galois group $G_{K}$ and formulate a conjecture on the structure of the graded Lie algebra associated to this filtration (Conjecture 2.8). This conjecture is an analogue of the so-called Deligne-Ihara conjecture in the case of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$. We then study some fundamental properties of the fixed field of $\ker(\rho_{X,p})$ of the pro-$p$ outer Galois representation associated to $X$, and state the main result.
The absolute Galois group $G_{K}$ inherits a descending central filtration $\{F^{m}G_{K}\}_{m\geq 1}$ by $F^{m}G_{K}\coloneqq\rho_{X,p}^{-1}(F^{m}\Gamma_{1,1})$. Note that we have $F^{1}G_{K}=G_{K(E[p^{\infty}])}$. We form graded quotients and their directed sum as
It is known that
  • Each $\mathfrak{g}_{m}$ is embedded into $F^{m}\Gamma_{1,1}/F^{m+1}\Gamma_{1,1}$, which is a free $\mathbb{Z}_{p}$-module of finite rank [NT93, Corollary (1.16), $(\mathrm{ii})$]. In particular, the $m$-th graded quotient $\mathfrak{g}_{m}$ is also a free $\mathbb{Z}_{p}$-module of finite rank.
  • Each $\mathfrak{g}_{m}\otimes\mathbb{Q}_{p}$ is isomorphic to a finite direct sum of (a finite direct sum of) $\mathbb{Q}_{p}(\boldsymbol{m})$’s as $G_{K}$-modules, where $\boldsymbol{m}\in\mathbb{Z}_{\geq 1}^{2}$ runs over indexes satisfying $|\boldsymbol{m}|=m$. This follows immediately from the $\mathrm{GL}_{2}(\mathbb{Z}_{p})$-equivariance of the commutative diagram given in [NT93, Theorem (1.14)].
We have the following lemma, relating the filtration with elliptic Soulé characters.
@@ 2.2. Analogue of the Deligne-Ihara conjecture and its consequence
    If $\kappa_{\boldsymbol{m}}\colon F^{1}G_{K}\to\mathbb{Z}_{p}(\boldsymbol{m})$ is nontrivial, then so is the character $\kappa_{\boldsymbol{m}}\mid_{F^{\lvert\boldsymbol{m}\rvert}G_{K}}$ restricted to ${F^{\lvert\boldsymbol{m}\rvert}G_{K}}$.
@@
In the following proof, we denote $\lvert\boldsymbol{m}\rvert$ by $m$, for simplicity. (1) Suppose that the restricted character $\kappa_{\boldsymbol{m}}\mid_{F^{m}G_{K}}$ is trivial. Then there exists an integer $1\leq n<m$ satisfying $\kappa_{\boldsymbol{m}}\mid_{F^{n+1}G_{K}}=0$ and $\kappa_{\boldsymbol{m}}\mid_{F^{n}G_{K}}\neq 0$. In particular, the $n$-th graded quotient $\mathfrak{g}_{n}\otimes\mathbb{Q}_{p}$ has a nontrivial $\chi^{\boldsymbol{m}}$-isotypic component, which is absurd since $n<m=\lvert\boldsymbol{m}\rvert$.
(2) It suffices to prove that the character $\kappa_{\boldsymbol{m}}$ vanishes when restricted to $F^{m+2}G_{K}$, since we have $F^{m+1}G_{K}=F^{m+2}G_{K}$ [Nak95, (4.2) Proposition]. By the construction of the power series $\alpha_{1,1}(\sigma)\in\mathbb{Z}_{p}[[T_{1},T_{2}]]$, it follows that, for every $\sigma\in F^{m+2}G_{K}$,

### v3

@@ 2.2. Elliptic analogues of Soulé characters
$\Gamma_{1,1}^{{\dagger}}\coloneqq\{f\in\operatorname{Aut}(\Pi_{1,1})\mid\text{$f$ preserves $\langle z\rangle$}\}.$
The weight filtration induces a filtration on $\Gamma^{{\dagger}}_{1,1}$ by taking the intersection
$F^{m}\Gamma^{{\dagger}}_{1,1}\coloneqq\Gamma^{{\dagger}}_{1,1}\cap F^{m}\tilde{\Gamma}_{1,1}.$
Observe that the normalizer subgroup of $\langle z\rangle$ in $\Pi_{1,1}$ coincides with $\langle z\rangle$ itself. It follows that the intersection $F^{m}\Gamma_{1,1}^{{\dagger}}\cap\operatorname{Inn}(\Pi_{1,1})$ is trivial for every $m\geq 3$ and otherwise coincides with $\langle\mathrm{inn}(z)\rangle$. Therefore, the natural projection induces an isomorphism
$F^{m}\Gamma_{1,1}^{{\dagger}}\to F^{m}\Gamma_{1,1}$
for every $m\geq 3$ and, otherwise, the kernel coincides with $\langle\mathrm{inn}(z)\rangle$ [Nak95, (4.4)]. These homomorphisms provide a group-theoretically natural way to lift elements of the pro-$p$ mapping class group.
We also define two subgroups $\Psi^{{\dagger}}_{1},\Psi^{{\dagger}}$ of the automorphism group of the maximal metabelian quotient of $\Pi_{1,1}$ by
@@
$\Psi^{{\dagger}}\coloneqq\{f\in\operatorname{Aut}\left(\Pi_{1,1}^{\mathrm{mab}}\right)\mid\text{$f$ preserves $\langle\bar{z}\rangle$}\}\quad\text{and}\quad\Psi^{{\dagger}}_{1}\coloneqq\operatorname{Ker}\left(\Psi^{{\dagger}}\to\operatorname{Aut}(\Pi^{\rm{ab}})\right),$
where $\bar{z}\in\Pi_{1,1}^{\mathrm{mab}}$ denotes the image of $z$. Note that each element $f\in\Psi^{{\dagger}}_{1}$ is uniquely determined by a pair $(f(x_{1})x_{1}^{-1},f(x_{2})x_{2}^{-1})$. Since $\Pi_{1,1}^{\mathrm{mab}}(2)$ is a free $\mathbb{Z}_{p}[[\Pi_{1,1}^{\mathrm{ab}}]]$-module generated by $\bar{z}$ [Iha86, Theorem 2], we can find a unique element $G_{i}(f)\in\mathbb{Z}_{p}[[\Pi_{1,1}^{\mathrm{ab}}]]$ satisfying
$f(x_{1})x_{1}^{-1}=G_{1}(f)z\quad\text{and}\quad f(x_{2})x_{2}^{-1}=G_{2}(f)z.$
In the following, we identify the completed group ring $\mathbb{Z}_{p}[[\Pi_{1,1}^{\rm{ab}}]]$ with the power series ring in two variables $\mathbb{Z}_{p}[[T_{1},T_{2}]]$ by $T_{i}\coloneqq x_{i}-1$ for $i=1,2$. We then obtain two power series from each element of $\Psi^{\dagger}_{1}$, and these two are related as follows:
@@
  1. (1)
    For $f\in\Psi^{{\dagger}}_{1}$, two power series $G_{1}(f)$ and $G_{2}(f)$ satisfy the relation
@@
  2. (2)
    If we write $H(f)\coloneqq\frac{G_{2}(f)}{T_{2}}=\frac{G_{1}(f)}{T_{1}}$, then we have an isomorphism
    $H\colon\Psi_{1}^{{\dagger}}\xrightarrow{\sim}\mathbb{Z}_{p}[[T_{1},T_{2}]];f\mapsto H(f).$
    In particular, the profinite group $\Psi_{1}^{{\dagger}}$ is abelian.
  3. (3)
    The $\operatorname{GL}_{2}(\mathbb{Z}_{p})$-action on $\Psi_{1}^{{\dagger}}$ induced by the exact sequence
    $1\to\Psi_{1}^{{\dagger}}\to\Psi\to\operatorname{GL}_{2}(\mathbb{Z}_{p})\to 1$
    makes an isomorphism $H\colon\Psi_{1}^{{\dagger}}\xrightarrow{\sim}\mathbb{Z}_{p}[[T_{1},T_{2}]](\mathrm{det})$ equivariant under the action of $\operatorname{GL}_{2}(\mathbb{Z}_{p})$. Here, $(\mathrm{det})$ denotes the twist by the determinant character.
@@
$s\colon F^{1}G_{K}=G_{K(E[p^{\infty}])}\to F^{3}\Gamma^{{\dagger}}_{1,1}$
@@
Hence the image of $F^{1}G_{K}$ under $\rho_{X,p}$ is contained in $F^{3}\Gamma_{1,1}$. By composing the inverse of the isomorphism $F^{3}\Gamma_{1,1}^{{\dagger}}\xrightarrow{\sim}F^{3}\Gamma_{1,1}$, we obtain the desired homomorphism $s\colon G_{K(E[p^{\infty}])}\to F^{3}\Gamma^{{\dagger}}_{1,1}$.
@@
We define a $\operatorname{Gal}(K(E[p^{\infty}])/K)$-equivariant homomorphism
$\alpha_{1,1}\colon G_{K(E[p^{\infty}])}^{\mathrm{ab}}\to\mathbb{Z}_{p}[[T_{1},T_{2}]](1)$
to be a compositum of
@@
    the isomorphism $H\colon\Psi^{{\dagger}}_{1}\xrightarrow{\sim}\mathbb{Z}_{p}[[T_{1},T_{2}]](\mathrm{det})$ constructed in Lemma 2.4.
In [Nak95], Nakamura obtained the explicit formula of the power series $\alpha_{1,1}$ for general once-punctured elliptic curves. We use the convention $0^{0}\coloneqq 1$ and regard $\mathbb{Z}_{p}[[T_{1},T_{2}]]$ as a subring of $\mathbb{Q}_{p}[[U_{1},U_{2}]]$ where $U_{i}\coloneqq\log(1+T_{i})$ in the following formula.
@@
for every $\sigma\in F^{1}G_{K}$, where $\kappa_{\boldsymbol{m}}\colon F^{1}G_{K}\to\mathbb{Z}_{p}$ is a character whose reduction modulo $p^{n}$ corresponds to the $p^{n}$-th roots of
@@
for every $n\geq 1$ via Kummer theory. Here $\theta(z,L)$ is the fundamental theta function [Nak95, (2.1)] and, $(\omega_{1,n})_{n},(\omega_{2,n})_{n}\in T_{p}(E)$ are defined in Lemma 2.1 (1).
Remark 2.6.1.
(1) Note that $\theta(z,L)$ is defined on $\mathbb{C}$ and is not periodic with respect to $L$. Hence we have to choose a lift of $a\omega_{1,n}+b\omega_{2,n}\in E(\mathbb{C})\cong\mathbb{C}/L$ to $\mathbb{C}$ to consider the value $\theta(a\omega_{1,n}+b\omega_{2,n},L)$. However, by [Nak95, (2.2) Proposition (1)], the values of $\theta(z,L)$ at any such two lifts coincide up to $p$-th power roots of unity. Since $K(E[p^{\infty}])$ contains all $p$-th power roots of unity, the Kummer characters associated to $p$-th power roots of $\theta(z,L)$ at any such two lifts also coincide. This resolves an ambiguity occurring in Theorem 2.6.
(2) Originally, the power series $\alpha_{1,1}$ and its explicit formula is defined and proved by using a certain basis of $\Pi_{1,1}$ coming from that of the topological fundamental group of $X(\mathbb{C})$. However, the proof of Theorem 2.6 given in [Nak95] works for our basis $\{x_{1},x_{2}\}$ as it is.
In the case of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$, it is known that coefficients of the universal power series for Jacobi sums correspond to the so-called Soulé characters $\{\chi_{m}\colon G_{\mathbb{Q}(\mu_{p^{\infty}})}^{\mathrm{ab}}\to\mathbb{Z}_{p}(m)\}_{m\geq 3,\mathrm{odd}}$ (see Ihara-Kaneko-Yukinari [IKY87, Theorem $\mathrm{A}_{2}$]). Essentially, these characters are introduced by Soulé [Sou81], and they enjoy arithmetically interesting properties. We refer the interested reader to the article of Ichimura-Sakaguchi [IS87], for example. Based on this analogy, we define their elliptic analogues as follows:
@@
For every $\boldsymbol{m}=(m_{1},m_{2})>\boldsymbol{1}$ such that $\lvert\boldsymbol{m}\rvert$ is even, we call the character $\kappa_{\boldsymbol{m}}$ the $\boldsymbol{m}$-th elliptic Soulé character.
Nakamura proved that some linear combinations of the elliptic Soulé characters are nontrivial [Nak95, (3.12)]. In our previous paper [Ish25], we conditionally proved that the elliptic Soulé characters arising from once-punctured CM elliptic curves are also nontrivial. We briefly summarize our previous results.
Since the power series $\alpha_{1,1}$ is compatible with the action of $\operatorname{Gal}(K(E[p^{\infty}])/K)$, the $\boldsymbol{m}$-th elliptic Soulé character $\kappa_{\boldsymbol{m}}$ is a $\operatorname{Gal}(K(E[p^{\infty}])/K)$-equivariant homomorphism
$\kappa_{\boldsymbol{m}}\colon G_{K(E[p^{\infty}])}^{\mathrm{ab}}\to\mathbb{Z}_{p}(\boldsymbol{m}).$
Recall that $\Omega$ is defined to be the maximal pro-$p$ extension of $K(p)$ unramified outside $p$.
@@
  1. (1)
    The $\boldsymbol{m}$-th elliptic Soulé character $\kappa_{\boldsymbol{m}}$ is trivial unless $\boldsymbol{m}$ is contained in $I$.
  2. (2)
    When $\boldsymbol{m}\in I$, the character $\kappa_{\boldsymbol{m}}$ factors through the maximal abelian quotient of $\operatorname{Gal}(\Omega/K(p^{\infty}))$. Moreover, it is nontrivial if $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(\boldsymbol{m}))$ is finite.
Here, we write
[…]
for the second étale cohomology group of the spectrum of the ring of $p$-integers of $K$ with coefficient in $\mathbb{Z}_{p}(\boldsymbol{m})$, whose finiteness is known to follow from a conjecture of Jannsen [Jan89, Conjecture 1]. It vanishes for every $\boldsymbol{m}\in(p-1)\mathbb{Z}_{\geq 1}^{2}$ by [Ish25, Lemma 4.7]. It is also finite when the coefficient is the $m$-th Tate twist $\mathbb{Z}_{p}(m)$ for some $m\geq 2$ by a classical result of Soulé [Sou79, page 287, Corollaire]. However, to the best of the author’s knowledge, the finiteness remains an open problem in general. We also have established a sufficient condition that elliptic Soulé characters are surjective, based on a relationship between elliptic Soulé characters and elliptic units:
@@
    Assume the class number of $K(p)$ is not divisible by $p$ and there are only two primes of $K(p)$ above $p$. Then $\kappa_{\boldsymbol{m}}$ is surjective for every $\boldsymbol{m}\in I$ such that $\boldsymbol{m}\not\equiv\boldsymbol{1}\bmod p-1$.
Although $\kappa_{\boldsymbol{m}}$ is not surjective if $\boldsymbol{m}\geq(2,2)$ and $\boldsymbol{m}\equiv\boldsymbol{1}\mod p-1$, we will see that it is nontrivial under the assumption of Theorem A, cf. Corollary A.2.
Now we discuss a relationship between Conjecture 2.3 with elliptic Soulé characters. One basic observation is the following:
@@
    If $\kappa_{\boldsymbol{m}}\colon F^{1}G_{K}\to\mathbb{Z}_{p}(\boldsymbol{m})$ is nontrivial, then so is its restriction to $F^{\lvert\boldsymbol{m}\rvert}G_{K}$}.
@@
We denote $\lvert\boldsymbol{m}\rvert$ by $m$ for simplicity. (1) Suppose that the restricted character $\kappa_{\boldsymbol{m}}\mid_{F^{m}G_{K}}$ is trivial. Then there exists an integer $1\leq n<m$ satisfying both $\kappa_{\boldsymbol{m}}\mid_{F^{n+1}G_{K}}=0$ and $\kappa_{\boldsymbol{m}}\mid_{F^{n}G_{K}}\neq 0$. In particular, the $n$-th graded quotient $\mathfrak{g}_{n}\otimes\mathbb{Q}_{p}$ has a nontrivial $\chi^{\boldsymbol{m}}$-isotypic component, which is absurd since $n$ is less than $m$.
(2) It suffices to prove that the character $\kappa_{\boldsymbol{m}}$ vanishes when restricted to $F^{m+2}G_{K}$, since we have $F^{m+1}G_{K}=F^{m+2}G_{K}$ by Proposition 2.2 (1). By the construction of the power series $\alpha_{1,1}(\sigma)\in\mathbb{Z}_{p}[[T_{1},T_{2}]]$, it follows that, for every $\sigma\in F^{m+2}G_{K}$,